# Approximating the coefficients of the Bessel functions

### v4

@@ 10.2. Upper bounds of the magnitudes of Bessel functions
In preparation for studying the uniform convergence of Bessel functions, we prove some upper bounds on the functions assuming that $\theta\in\theta(\mathcal{R})$ is nonnegative so that we can apply Theorem˜1.4.
@@
First, observe that from Theorem˜1.4,
$J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}(\epsilon)$
for a measure $\mu_{a}$ over $\mathbb{R}^{N}$ which is supported on the convex hull of $H(\mathcal{R})a$ and is invariant with respect to the action of $H(\mathcal{R})$. In particular,
$\left\lVert J^{\mathcal{R}(\theta)}_{a(N)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}x_{i}h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon).$
@@
(A): We use Theorem˜1.4. Note that for $z\in\mathbb{C}^{N}$, we let $\text{Re}(z)$ denote $(\text{Re}(z_{i}))_{i\in[N]}$ and for $r\in\mathbb{R}^{N}$, we let $|r|$ denote $(|r_{i}|)_{i\in[N]}$. First, observe that
$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}e^{\left\langle\text{Re}(x),\epsilon\right\rangle}d\mu_{a}(\epsilon)\leq\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon).$
Next, note that since $\mu_{a}$ is permutation invariant because $H(\mathcal{R})$ contains the symmetric group,
$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)=\int_{\mathbb{R}^{N}}1+\sum_{\nu\in\Gamma}\frac{M_{\nu}(|\text{Re}(x)|)M_{\nu}(|\epsilon|)}{M_{\nu}(1,\ldots,1)\nu!}d\mu_{a}(\epsilon).$
@@
(31) $\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\nu!}=e^{R\sum_{i=1}^{N}|\text{Re}(x)_{i}|}.$
@@
$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}|\nu|!R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\pi(\nu)}=\sum_{d=0}^{\infty}\left(R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d},$
@@
10.3. The uniform convergence of Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime
@@
    $m_{2d}=\sum_{\pi\in NC^{\text{even}}(2d)}W^{BC}(\pi)(2c_{0},2c_{1})\prod_{B\in\pi}|B|c_{(|B|)}$
@@
(A): By Lemma˜10.5, $\zeta\triangleq\{J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
@@
Afterwards, by part (B) of Lemma˜10.7, we have that $\zeta$ converges to $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$ uniformly over compact subsets of $\Omega$.
(B) and (C): We can use the same method as (A); to replace Corollary˜9.4, we use Corollary˜9.8 for (B) and Corollary˜9.14 for (C). For (C), observe that the terms with all odd degrees are eliminated over $\Omega$. ∎
@@ 10.4. The uniform convergence of Bessel functions for Vershik-Kerov
In this subsection, we reprove the results of the papers [assiotis, rank_infinity] and extend these results to the $D^{N}$ root system. The setting that we consider is equivalent to the setting of Vershik-Kerov sequences that the papers consider, see Remark˜10.13. First, we prove the analogues of Theorems˜6.10, 7.4 and 8.21.
@@
See Theorems˜8.10 and 8.16. ∎
Theorem˜10.12 is an example of an application of the results of this paper. Part (A) of the theorem appears in [assiotis] while parts (A) and (B) appear in [rank_infinity]. Part (C) has not appeared previously, although it is straightforward to deduce. See Remark˜10.13 for discussion on how the setting of the theorem is equivalent to the setting of Vershik-Kerov sequences.
@@
(A): We include the proof of this result from [rank_infinity] for completeness. Using Theorem˜1.4 and Hölder’s inequality gives that
@@
Afterwards, using Lemma˜10.6 gives that
@@
It is then is clear that $\zeta\triangleq\{J_{a(N)}^{A^{N-1}(\theta)}(x_{1},\ldots,x_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{rM}\,\forall i\in[r]\}$. Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$; the application of Corollary˜9.2 is the difference between this proof and that of [rank_infinity]. We can then apply part (B) of Lemma˜10.7 to conclude the proof.
(B) and (C): We follow the same method as the proof of (A). For (B), we use part (B) of Lemma˜10.6 and Corollary˜9.6 and for (C), we use part (C) of Lemma˜10.6 and Corollary˜9.11. ∎
@@
From [assiotis]*Proposition 2.3, the condition that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N^{d}}=m_{d}$ for all $d\geq 1$ implies that $\{a(N)\}_{N\geq 1}$ is a Vershik-Kerov sequence after it is reordered. Then, from [rank_infinity], we can set $M=\alpha+\epsilon$ in Theorem˜10.12 for any $\epsilon>0$, where $\alpha=\lim_{N\rightarrow\infty}\max_{i\in[N]}\left|\frac{a(N)_{i}}{N}\right|$. This will imply uniform convergence over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{r\alpha}\}$, which is the version of the result that appears in [assiotis, rank_infinity].
@@ 10.5. Bessel generating functions for exponentially decaying probability measures
Rather than consider the Bessel function $J_{a}^{\mathcal{R}(\theta)}(x)$ for a single value of $a$, we can consider the average of $J_{a}^{\mathcal{R}(\theta)}(x)$ over a distribution of $a$. The resulting function is referred to as the Bessel generating function and has been studied previously in [gaussianfluctuation, matrix, limits_prob_measures, rectangularmatrix]. This notion is also related to the Dunkl transform discussed in [dunklbound].
Suppose $\mu$ is a Borel probability distribution over $\mathbb{C}^{N}$. Then, the Bessel generating function associated to $\mu$ and $\mathcal{R}(\theta)$ is
$G_{\mu}^{\mathcal{R}(\theta)}(x)\triangleq\operatorname{\mathbb{E}}_{a\sim\mu}[J_{a}^{\mathcal{R}(\theta)}(x)].$
However, we must place restrictions on $\mu$ so that the Bessel generating function converges. In particular, we consider the class of exponentially decaying probability measures studied in [limits_prob_measures], which is a modification of the class of measures studied in [matrix].
@@
converges because $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\left\lVert a\right\rVert_{2}\left\lVert x\right\rVert_{2})$ and $\mu$ exponentially decays at rate $R$. Furthermore, we can show that the integral is holomorphic over $B_{R}(0)$ by following [matrix]*Proof of Lemma 2.9. Note that we have just proved the nonsymmetric version of (A), which can be used to prove (A).
@@
The papers [matrix, limits_prob_measures] only consider the $A^{N}$ root system. However, (A) and (D) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound]. Furthermore, observe that (A) implies (B). However, we require the exponentially decaying condition to prove (C) and (D).
Using Lemma˜10.15 and part (A) of Theorem˜1.1, we are able to deduce the following corollary for the $A^{N-1}$ root system. It is also straightforward to deduce the analogous results for the $BC^{N}$ and $D^{N}$ root systems. The corollary generalizes the results of [limits_prob_measures].
@@
For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{N}\}$, define the coefficients $c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ and for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{N}\}$, define the coefficients $c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ as they are defined in part (B) of Lemma˜10.15. Then, the following are equivalent.
@@
  2. (a’)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
Condition (b) of Corollary˜10.17 has been studied in previous works such as [llnclt, matrix, limits_prob_measures, rectangularmatrix, jackgenfunc]. When the condition is satisfied, the sequence of probability measures satisfies a law of large numbers. Next, we state the analogous result for the $BC^{N}$ root system.
@@
Suppose $\mu_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{C}^{N}$ for all $N\geq 1$. Assume that $\theta_{0},\theta_{1}\in\mathbb{C}$ have nonnegative real part for all $N\geq 1$, $\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$, define the coefficients $c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ and for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{\text{even};\,N}\}$, define the coefficients $c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{\text{even};\,N}$ as they are defined in part (B) of Lemma˜10.15. Then, the following are equivalent.
@@
  2. (a’)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
For the type D analogue of Corollary˜10.17, we adjust the notation used to match the notation of part (C) of Theorem˜1.1. Therefore, we first prove the following lemma, which is analogous to part (B) of Lemma˜10.15.
@@
$G_{\mu}^{D^{N}(\theta)}=\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{\mathcal{R}(\theta)}(\mu)b\right)+ed_{0}^{D^{N}(\theta)}(\mu)+e\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{\mathcal{R}(\theta)}(\mu)b\right)$
@@
Define the coefficient $d_{0}^{D^{N}(\theta)}(\mu)$, the coefficients $c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N}),d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ for $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$, and the coefficients $c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N}),d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{\text{even};\,N}$ for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{\text{even};\,N}\}$ as they are defined in Lemma˜10.19. Consider the following statements.
@@
  2. (a’)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even};\,N}$, $\lim_{N\rightarrow\infty}\\ \frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  4. (b’)
    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even};\,N}$, $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.

### v5

@@ 10.2. Upper bounds of the magnitudes of Bessel functions
In preparation for studying the uniform convergence of Bessel functions, we prove various upper bounds on them.
Lemma 10.4.
Suppose $\theta\in\Theta(\mathcal{R})$ and $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$. Then,
$|J_{a}^{\mathcal{R}(\theta)}(x)|\leq\exp\left(\max_{h\in H(\mathcal{R})}\text{Re}(\left\langle ha,x\right\rangle)\right).$
Proof.
By [dunklbound]*Theorem 3.1,
$\left(\frac{1}{|H(\mathcal{R})|}\sum_{h\in H(\mathcal{R})}E_{ha}^{\mathcal{R}(\theta)}(x)\right)^{2}\leq\frac{1}{|H(\mathcal{R})|}\sum_{h\in H(\mathcal{R})}E_{ha}^{\mathcal{R}(\theta)}(x)^{2}\leq\exp\left(2\max_{h\in H(\mathcal{R})}\text{Re}(\left\langle ha,x\right\rangle)\right),$
which finishes the proof. ∎
@@
First, observe that from Theorem˜1.3,
$J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)$
for a measure $\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}$ over $\mathbb{R}^{N}$ which is supported over the convex hull of $H(\mathcal{R})a$ and is invariant with respect to the action of $H(\mathcal{R})$. In particular,
$\left\lVert J^{\mathcal{R}(\theta)}_{a}(x)\right\rVert_{2}$ $=\left\lVert\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}x_{i}h^{\prime}\epsilon_{i}}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)\right\rVert_{2}$
$\leq\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}\epsilon_{i}}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon).$
@@
(A): We use Theorem˜1.3. Note that for $z\in\mathbb{C}^{N}$, we let $\text{Re}(z)$ denote $(\text{Re}(z_{i}))_{i\in[N]}$ and for $r\in\mathbb{R}^{N}$, we let $|r|$ denote $(|r_{i}|)_{i\in[N]}$. First, observe that
$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}$ $=\left\lVert\int_{\mathbb{R}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}e^{\left\langle\text{Re}(x),\epsilon\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)$
$\leq\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon).$
Next, note that since $\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}$ is permutation invariant because $H(\mathcal{R})$ contains the symmetric group,
$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)=\int_{\mathbb{R}^{N}}1+\sum_{\nu\in\Gamma}\frac{M_{\nu}(|\text{Re}(x)|)M_{\nu}(|\epsilon|)}{M_{\nu}(1,\ldots,1)\nu!}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon).$
@@
(22) $\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\nu!}=e^{R\sum_{i=1}^{N}|\text{Re}(x)_{i}|}.$
@@
$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\nu_{a}^{\text{sym; $\mathcal{R}(\theta)$}}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}|\nu|!R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\pi(\nu)}=\sum_{d=0}^{\infty}\left(R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d},$
@@
10.3. The uniform convergence of Bessel functions in the high temperature regime
@@
    $m_{2d}=\sum_{\pi\in NC^{\text{even}}(2d)}W^{BC}(\pi)(c_{0},c_{1})\prod_{B\in\pi}|B|c_{(|B|)}$
@@
(A): By Lemma˜10.6, $\zeta\triangleq\{J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
@@
Afterwards, by part (B) of Lemma˜10.8, we have that $\zeta$ converges to $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$ uniformly over compact subsets of $\Omega$.
(B) and (C): We can use the same method as (A); to replace Corollary˜9.4, we use Corollary˜9.8 for (B) and (20) and Corollary˜9.8 for (C). For (C), observe that the terms with all odd degrees are eliminated over $\Omega$. ∎
@@ 10.4. The uniform convergence of Bessel functions for Vershik-Kerov
In this subsection, we reprove the results of the papers [assiotis, rank_infinity] and extend these results to the $D^{N}$ root system. The setting that we consider is equivalent to the setting of Vershik-Kerov sequences that the papers consider, see Remark˜10.14. First, we prove the analogues of Theorems˜6.10, 7.4 and 8.9.
@@
See Theorem˜7.1 with $c=0$ and Lemma˜8.8. ∎
Theorem˜10.13 is an example of an application of the results of this paper. Part (A) of the theorem appears in [assiotis] while parts (A) and (B) appear in [rank_infinity]. Part (C) has not appeared previously, although it is straightforward to deduce. See Remark˜10.14 for discussion on how the setting of the theorem is equivalent to the setting of Vershik-Kerov sequences.
@@
(A): We include the proof of this result from [rank_infinity] for completeness. Using Theorem˜1.3 and Hölder’s inequality gives that
@@
Afterwards, using Lemma˜10.7 gives that
@@
It is then is clear that $\zeta\triangleq\{J_{a(N)}^{A^{N-1}(\theta)}(x_{1},\ldots,x_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{rM}\,\forall i\in[r]\}$. Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$; the application of Corollary˜9.2 is the difference between this proof and that of [rank_infinity]. We can then apply part (B) of Lemma˜10.8 to conclude the proof.
(B) and (C): We follow the same method as the proof of (A). For (B), we use part (B) of Lemma˜10.7 and Corollary˜9.6 and for (C), we use part (C) of Lemma˜10.7, (20), and Corollary˜9.6. ∎
@@
From [assiotis]*Proposition 2.3, the condition that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N^{d}}=m_{d}$ for all $d\geq 1$ implies that $\{a(N)\}_{N\geq 1}$ is a Vershik-Kerov sequence after it is reordered. Then, from [rank_infinity], we can set $M=\alpha+\epsilon$ in Theorem˜10.13 for any $\epsilon>0$, where $\alpha=\lim_{N\rightarrow\infty}\max_{i\in[N]}\left|\frac{a(N)_{i}}{N}\right|$. This will imply uniform convergence over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{r\alpha}\}$, which is the version of the result that appears in [assiotis, rank_infinity].
@@ 10.5. Bessel generating functions for exponentially decaying probability measures
Recall that the Bessel generating function is defined in Definition˜1.1. We must place restrictions on $\mu$ so that the expected value converges. In particular, we consider the class of exponentially decaying probability measures studied in [limits_prob_measures], which is a modification of the class of measures studied in [matrix].
@@
converges because $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\left\lVert a\right\rVert_{2}\left\lVert x\right\rVert_{2})$ and $\mu$ exponentially decays at rate $R$. Furthermore, we can show that the integral is holomorphic over $B_{R}(0)$ by following [matrix]*Proof of Lemma 2.9. Note that we have proved the nonsymmetric version of (A), which can be used to prove (A).
@@
The papers [matrix, limits_prob_measures] only consider the $A^{N-1}$ root system. However, (A) and (D) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound]. Furthermore, observe that (A) implies (B). However, we require the exponentially decaying condition to prove (C) and (D).
Using Lemma˜10.16 and part (A) of Theorem˜1.2, we are able to deduce the following corollary for the $A^{N-1}$ root system. It is also straightforward to deduce the analogous results for the $BC^{N}$ and $D^{N}$ root systems. The corollary generalizes the results of [limits_prob_measures].
@@
For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{N}\}$, define the coefficients $c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ as they are defined in part (B) of Lemma˜10.16. Then, the following are equivalent.
@@
Condition (b) of Corollary˜10.18 has been studied in previous works such as [llnclt, matrix, limits_prob_measures, rectangularmatrix, jackgenfunc]. When the condition is satisfied, the sequence of probability measures satisfies a law of large numbers. Next, we state the analogous result for the $BC^{N}$ root system.
@@
[…]
For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$, define the coefficients $c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ as they are defined in part (B) of Lemma˜10.16. Then, the following are equivalent.
@@
For the type D analogue of Corollary˜10.18, we adjust the notation used to match the notation of part (C) of Theorem˜1.2. Therefore, we first prove the following lemma, which is analogous to part (B) of Lemma˜10.16.
@@
$G_{\mu}^{D^{N}(\theta)}=\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{D^{N}(\theta)}(\mu)b\right)+ed_{0}^{D^{N}(\theta)}(\mu)+e\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{D^{N}(\theta)}(\mu)b\right)$
@@
Define the coefficients $d_{0}^{D^{N}(\theta)}(\mu)$ and $c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N}),d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ as they are defined in Lemma˜10.20 for $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$. Consider the following statements.
@@
    It is the case that $\lim_{N\rightarrow\infty}d_{0}^{D^{N}(\theta)}(\mu_{N})=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even};\,N}$, $\lim_{N\rightarrow\infty}\\ \frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
Then, (a) and (c) are equivalent and (b) and (d) are equivalent.
Remark 10.22.
It is straightforward to establish analogues of Corollaries˜10.18, 10.19 and 10.21 in the $\theta N\rightarrow c\in\mathbb{C}$ regime.